A newly discovered pollutant is fossil fuels was characterized to determine its composition. Elemental analysis shows the presence of carbon,hydrogen and sulfur elements in the compound. When burned with excess oxygen 0.397 g carbon dioxide 0.190g water and 0.193 g sulfur dioxide were found as products. What is the mass percent of the pollutant in fuel??


1
Expert's answer
2021-11-19T08:42:01-0500

0.397+0.19+0.193= 0.78

0.19/0.78=0.244×100

= 24.4%